OpenStudy (momo):

What is the slope of a line perpendicular to the line 4x - 5y = 8?

OpenStudy (anonymous):

y = (8 - 4x) / 5 y = 8/5 - (4/5)x y = -(4/5)x + 8/5 the perp. slope is -1 / (-4 / 5) = 5 / 4

OpenStudy (momo):

i was thinking that. so it would be -5/4?

OpenStudy (anonymous):

if M is a slope, then its perpendicular is -1 / M in this case, M = -4/5 so perp. to M is 5 / 4

OpenStudy (momo):

well, that isn't one of my options. my options are 8, -4, 4/5, and -5/4 i was thinking maybe the - would be cancelled out and it would be 4/5

OpenStudy (anonymous):

y = Mx + b y = (4/5)x - (8/5) perp = -1/M = -1/ (4/5) = -5/4

OpenStudy (momo):

awesome! thanks :). my original thought haha

OpenStudy (anonymous):

np (: